plzzsolve this
After a discount of $259.75, a distributor paid $892.75 to purchase an item from a manufacturer. What rate of discount was applied to the purchase? (Round your answer to 1 decimal place.)

Answer by rothauserc(4718)

Let x be the total amount and r be the rate, then
x - 259.75 = 892.75
x = 892.75 + 259.75 = 1152.50
:
1152.50r = 259.75
r = 259.75 / 1152.50 = 0.2253
Rate of discount is 22.5%